# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

### v2

@@ 3.3 Experimental Settings
Each regressor is evaluated using RMSE as our performance metric and evaluated 5 times. We set our experiment settings as follows:
@@
  • Unseen: The performance records can be divided into two categories: (1) records with “seen" languages and (2) records with “unseen" languages. “Unseen" languages refer to languages that are not present in the pre-training LM data, while “seen" languages denote those that are present. In this setting, the regressor is trained using records of “seen" languages and tested using records of “unseen" languages.
  • Cross-Dataset: We train the regressor using performance records from the English-centric dataset and test it using records from the Many-to-Many Languages dataset. We opt not to reverse this setup as the dataset exhibits no domain shift and contains fewer performance records.
@@ 4 Results and Analysis
4.1 English-centric Results
Table 1 shows the overall results on the English-centric dataset. ProxyLM remarkably outperforms all existing baselines. We find that incorporating all proxy models (Ensemble) is the most effective for prediction, leading to a 2.29$\times$ averaged reduction in RMSE across all experimental settings compared to the best baseline. We observe that using the No FT estimated model to predict the performance of their fine-tuned models is surprisingly useful in all settings, especially for NLLB, where the model already has decent machine translation quality on LRLs. This observation is supported by our findings within the XGBoost model that the NLLB No FT feature has the highest importance score among all features, as shown in Figure 16 in the Appendix. Further, using SMaLL-100 fine-tuned performance provides useful estimations for settings involving M2M100 as the estimated model. This may indicate that the performance of a model with similar architecture can be a good estimator for the performance of the larger estimated model. In other words, the choice of proxy model to help prediction matters. Feature importance analysis from the XGBoost model supports this, revealing that the SMaLL-100 fine-tuned feature has the highest importance score among all features, as shown in Figure 15 in the Appendix.
Our analysis also indicates that XGBoost outperforms other regression models, across all evaluated settings. Both XGBoost and LGBM, which are gradient-boosting and tree-based learning methods, demonstrate superior performance metrics across all settings. Their robustness and efficiency as non-linear models are evident when compared to linear models, such as Poly2 and Poly3. Poly2 and Poly3 regressors, which employ second-degree and third-degree polynomial regression approaches respectively, tend to generate lower scores. This diminished performance is largely attributed to their limitations in capturing the nonlinear relationships inherent in the data, leading to suboptimal results. We further present the results by the language vitality in Table 2. The overall difference in RMSE between LRLs and medium-resource languages (MRLs) is relatively small, except when using SMaLL-100 as proxy models. An interesting observation here on SMaLL-100 in predicting the NLLB model is that the No FT model can predict LRLs much better than the fine-tuned counterpart, and the fine-tuned model better predicts MRLs.
[…]
NLPerf Xia et al. (2020a) with different regressors
XGBoost 9.97 $\pm$ 0.35 10.52 $\pm$ 0.73
Poly2  Khiu et al. (2024) 9.27 $\pm$ 0.00 21.43 $\pm$ 0.00
Poly3  Khiu et al. (2024) 8.58 $\pm$ 0.00 19.81 $\pm$ 0.00
LGBM  Ye et al. (2021a) 9.24 $\pm$ 0.24 15.19 $\pm$ 1.00
@@ 4.1 English-centric Results
Transformer 6.65 $\pm$ 0.26 10.69 $\pm$ 0.90
SMaLL-100 5.53 $\pm$ 0.98 7.68 $\pm$ 2.11
SMaLL-100 (No FT) 9.08 $\pm$ 0.63 16.66 $\pm$ 1.08
Estimated Model (No FT) 9.45 $\pm$ 0.78 7.10 $\pm$ 0.69
Ensemble 4.30 $\pm$ 0.17 6.21 $\pm$ 0.50
Table 4: Cross-Dataset test results in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning". The reported results are experiments using the LGBM regressor. Ensemble denotes combining all four proxy models.
4.2 Many-to-Many Languages Results
@@
4.3 Cross-Dataset Results
Table 4 illustrates model performance in the Cross-Dataset setup, showcasing the superior performance of ProxyLM with LGBM. We use LGBM instead of XGBoost because LGBM performs better. The results highlight that ProxyLM with Ensemble significantly reduces RMSE compared to the best baseline by 2$\times$ and 1.69$\times$ for M2M100 and NLLB, respectively. This displays consistent performance across datasets and languages that are not encountered during the regressor’s training, including unseen languages for the pre-trained LMs. However, the No FT models exhibit inconsistency compared to other proxy models. We also find varied performance between M2M100 and NLLB, which may be attributed to the English-centric dataset containing only seen languages for NLLB and lacking unseen language examples for the regressor. This highlights the significance of incorporating unseen language instances in training for more robust predictions.
@@ 4.3 Cross-Dataset Results
Figure 3: RMSE decreases as XGBoost’s training set size increases. Training set size refers to the number of English-centric past performance records.
4.4 Ablation Study
Figure 2 highlights the impact of features used in ProxyLM in the LOLO setting with XGBoost. Utilizing proxy models as features leads to a significant reduction in RMSE across all scenarios, showcasing their importance compared to other features. For the English-centric dataset, including language and dataset features alongside proxy models enhances performance. Dataset features alone show better improvement than language features alone, but the combination of both yields the best performance. On the other hand, for the Many-to-Many Languages dataset, the benefits of incorporating dataset and language features are less pronounced, especially for the M2M100 model, and there may even be a performance dip for the NLLB model due to the dataset’s lack of domain shift.
4.5 Diminishing Returns with Increasing Training Set Size
In Figure 3, we examine the training of the XGBoost regressor using different numbers of English-centric past performance records as the training dataset. While the regressor’s performance shows enhancement with an expanding training size, the incremental benefits start diminishing once the training set surpasses about 400 past performance records. This observation implies that, across datasets, there exists a threshold where the advantages of incorporating additional past performance records begin to exhibit diminishing returns.
@@ 4.5 Diminishing Returns with Increasing Training Set Size
M2M100 338 s 3.94 hrs (4.77$\times$) 1.42 hrs (7.32$\times$)
NLLB 492 s 9.4 hrs (11.48$\times$) 7.21 hrs (37.08$\times$)
@@
SMaLL-100 201 s 2.5 hrs (3.03$\times$) 1.03 hrs (5.29$\times$)
Transformer 60 s 0.83 hrs (1$\times$) 0.19 hrs (1$\times$)
Table 5: Comparison of LMs’ inference time (in seconds) and fine-tuning time (in hours) for one experimental run. The multiplier of fine-tuning time is relative to the Transformer model. All times were calculated using the interquartile mean to ignore outliers. Since the inference time for English-centric and Many-to-Many Languages datasets are roughly the same, only one column is provided.
Regressors English-centric Many-to-Many Langs.
XGBoost 129.29 s 69.81 s
Poly2 0.06 s 0.03 s
Poly3 0.07 s 0.06 s
LGBM 702.12 s 173.17 s
MF N/A 465.37 s
Table 6: Regressor models training time comparison in seconds (s).
Figure 4: Detailed results of XGBoost with ProxyLM Ensemble on M2M100 model under the LOLO setting using the English-centric dataset from Table 1. The results are grouped by (a) Joshi Class and (b) language family that follows the mapping which is provided in Appendix A.1; (c) shows the scatter plot illustrating the correlation of spBLEU scores between the ProxyLM’s prediction and estimated LM, with the light gray dashed line representing the line of equality ($y=x$) with $R^{2}=0.88$ and black dashed line representing Locally Weighted Scatterplot Smoothing (LOWESS) curve to represent the trend.
@@ 4.6 Time Efficiency
Table 5 compares the fine-tuning and inference times required for the estimated and proxy models. The results demonstrate that fine-tuning proxy models or direct inference from any model is remarkably faster than fine-tuning all estimated models. Table 6 further illustrates this point, showing only a minimal trade-off in the time needed to train the regressor models. This additional training time is relatively negligible, highlighting the efficiency of using proxy models.
@@ 4.7 Performance by Language Categories
In Figure 4, we present detailed XGBoost results with ProxyLM Ensemble on the M2M100 model under the English-centric LOLO experiment, grouped by language categories. Based on the Locally Weighted Scatterplot Smoothing (LOWESS)  Cleveland (1979) curve depicted in Figure 4(c), our method consistently maintains unbiased predictions for spBLEU scores below 40 across various language types. However, as the spBLEU score increases, the availability of data points diminishes, leading to our method under-predicting the performance compared to the true spBLEU score. Outliers observed in Kartvelian languages and Indo-European languages with Joshi class 3 may have contributed to this discrepancy in prediction. These observations suggest that increasing the number of data points covering higher spBLEU scores may help mitigate the bias in prediction. Further experiment details are available in Appendix B.
@@ 5 Related Work
The prediction performance of machine learning algorithms has been mainly explored in two research directions: (1) predict the model performance during the training runtime, and (2) predict the model performance by providing extracted features from the dataset Xia et al. (2020a).
Performance Prediction During the Training Runtime
@@
Performance Prediction Using Extracted Features
The latter aims to predict the model performance by learning a correlation between input features and the final evaluation metric. Birch et al. (2008a) identify strong predictive features such as the amount of reordering, the morphological complexity of the target language, and the historical relatedness of the two languages. Xia et al. (2020a) leverage extracted dataset features and typological database language representations. Ye et al. (2021a) introduce the use of confidence intervals and calibration with various regressor algorithms for reliable performance prediction. Schram et al. (2023) apply Bayesian matrix factorization for performance prediction on multilingual NLP tasks. In this work, we focus on exploring the latter. Existing approaches have shown promise using linear regression and gradient-boosting trees Birch et al. (2008b); Xia et al. (2020b); Srinivasan et al. (2021); Ye et al. (2021b). These studies have considered data size, typological features, and language similarity as factors contributing to the model performance.
@@ 6 Conclusion
In this paper, we introduce ProxyLM, a novel framework designed to predict the performance of LMs by leveraging proxy models including for LRLs. By utilizing proxy models as substitutes to estimate the performance of the target model, we strategically employ smaller LMs and off-the-shelf models without additional fine-tuning. This framework is highly scalable to multiple proxy models and is task-agnostic, making it applicable to a wide range of downstream NLP tasks. Our streamlined approach showcases substantial advancements in prediction accuracy compared to standard baselines and exhibits strong generalization capabilities across varied scenarios.
@@
This paper focuses exclusively on two estimated models: M2M100 and NLLB, to evaluate our proposed framework. For demonstration purposes, we concentrate on the usage of specific models, namely the Transformer model, the SMaLL-100 model, and the No FT models, to illustrate the effectiveness of our proxy models. The M2M100 and NLLB models were selected due to their prominence and relevance in the field of multilingual translation tasks. These models serve as robust benchmarks for assessing the performance and reliability of our proxy-based framework. By using these well-regarded models, we aim to provide compelling evidence of the capabilities and advantages of ProxyLM. While our proposed framework is evaluated solely within the context of machine translation, it is not confined to this application alone. The framework is designed to be versatile and can be extended to a variety of other downstream tasks. We plan to explore these additional applications in future work.
Some other possible avenues for future work could involve a deeper investigation into which proxy models are more effective for enhancing performance prediction in specific settings. Our findings suggest that one proxy model can outperform another in different scenarios, making it crucial to carefully select the most relevant proxy models to maximize the benefits of our approach. Additionally, developing methodologies for collecting relevant past performance records could provide better insights and improve the generalization and accuracy of our framework. Past performance records may provide better information gain than others, potentially minimizing the number of performance records required for a more robust and accurate predictor.

### v3

@@ 3.5 Experimental Settings
Each regressor is evaluated using RMSE as our performance metric and evaluated 5 times. For all tasks, we set our experiment settings as follows:
@@
To test the robustness of ProxyLM, we also provide two additional setups specifically for MT:
  • Unseen: The performance records for MT can be divided into two categories: (1) records with “seen" languages and (2) records with “unseen" languages. “Unseen" languages refer to languages that are not present in the pre-training LM data, while “seen" languages denote those that are present. In this setting, the regressor is trained using records of “seen" languages and tested using records of “unseen" languages.
  • Cross-Dataset: The regressor can be trained using performance records from the English-centric dataset and tested using the Many-to-Many Languages dataset. We opt not to reverse this setup as the Many-to-Many dataset exhibits no domain shift and contains fewer performance records.
@@ 4 Results and Analysis
4.1 Machine Translation
4.1.1 English-centric Results
Table 1 shows the overall results on the English-centric dataset using spBLEU. ProxyLM remarkably outperforms all existing baselines. We find that incorporating all proxy models (Ensemble) is the most effective for prediction, leading to a 2.41$\times$ averaged reduction in RMSE across all experimental settings compared to the best baseline. Note that the significant improvement remains consistent when evaluated using a different metric, such as COMET-22, which yields a 2.00× averaged reduction in RMSE across experimental settings compared to the best baseline, as shown in Table 14 in Appendix Section A. We observe that using the No FT estimated model to predict the performance of their fine-tuned models is surprisingly useful in all settings, especially for NLLB, where the model already has decent machine translation quality on LRLs. This observation is supported by our findings within the XGBoost model that the NLLB No FT feature has the highest importance score among all features, as shown in Figure 9 in the Appendix. Furthermore, using SMaLL-100 fine-tuned performance provides useful estimations for settings involving M2M100 as the estimated model. This may indicate that the performance of a model with similar architecture can be a good estimator for the performance of the larger estimated model. In other words, the choice of proxy model to help prediction matters. Feature importance analysis from the XGBoost model supports this, revealing that the SMaLL-100 fine-tuned feature has the highest importance score among all features, as shown in Figure 8 in the Appendix.
Intent Classification Slot Filling
Models Random LOLO Avg. Random LOLO Avg.
Aya23$\downarrow$ LLaMA3$\downarrow$ Aya23$\downarrow$ LLaMA3$\downarrow$ Aya23$\downarrow$ LLaMA3$\downarrow$ Aya23$\downarrow$ LLaMA3$\downarrow$
NLPerf Xia et al. (2020) with different regressors
XGBoost 0.0761 0.0191 0.1573 0.0581 0.0777 0.0693 0.0548 0.1219 0.1093 0.0888
Poly2 Khiu et al. (2024) 0.1996 0.0979 0.2075 0.0918 0.1492 0.1396 0.1412 0.1418 0.1414 0.1410
Poly3 Khiu et al. (2024) 0.1990 0.0969 0.2191 0.0925 0.1519 0.1393 0.1401 0.1448 0.1413 0.1414
LGBM Ye et al. (2021) 0.0839 0.0198 0.1545 0.0558 0.0785 0.0692 0.0557 0.1218 0.1152 0.0905
@@ 4.1.1 English-centric Results
SmolLM (135M) 0.0676 0.0171 0.1273 0.0455 0.0644 0.0618 0.0538 0.1004 0.0953 0.0778
SmolLM (360M) 0.0604 0.0157 0.1118 0.0441 0.0580 0.0562 0.0506 0.0844 0.0868 0.0695
BLOOMZ (560M) 0.0692 0.0179 0.1283 0.0482 0.0659 0.0618 0.0540 0.1023 0.0995 0.0794
Ensemble 0.0609 0.0164 0.1112 0.0442 0.0582 0.0561 0.0508 0.0830 0.0884 0.0696
Table 2: Intent classification and slot filling results using accuracy and micro-F1 score, respectively, with average RMSE (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. Avg represents the average of the results across the row for each respective task. The reported results use XGBoost as the regressor for both intent classification and slot filling. Ensemble denotes combining all three proxy models. The detailed breakdown of this result with the standard deviation can be seen in the Appendix Section C.
4.1.2 Many-to-Many Languages Results
@@
4.2 Intent Classification and Slot Filling
Table 2 presents the overall results for both intent classification and slot filling tasks. In these experiments, ProxyLM employs XGBoost as the regressor for intent classification and LGBM for slot filling, as LGBM exhibited superior performance during training cross-validation for the slot filling task. Consistent with our findings in MT, ProxyLM outperforms all existing baselines in both tasks. Notably, the SmolLM (360M) model emerges as the most effective proxy model, achieving an average RMSE reduction of 1.34$\times$ in intent classification and 1.28$\times$ in slot filling across all experimental settings when compared to the best baseline. As with MT, the choice of proxy model significantly impacts prediction performance. Based on the feature importance analysis in Figure 10-13 in the Appendix, SmolLM (360M) exhibits the highest importance score among all proxy models in both tasks. This suggests that the size of the proxy models may not be a reliable indicator of their effectiveness.
5 Ablation Study
5.1 Robustness of ProxyLM
Figure 2 illustrates the performance of our models in both the Unseen and Cross-Dataset setups, highlighting the robustness results achieved by ProxyLM. For the Cross-Dataset evaluation, we opted for LGBM instead of XGBoost, as it demonstrated better performance during cross-validation on the training set. ProxyLM with Ensemble shows a significant reduction in RMSE compared to the best baseline: a 1.84$\times$ reduction in the Unseen setup, and reductions of 2.15$\times$ and 1.78$\times$ for M2M100 and NLLB in the Cross-Dataset scenario, respectively. This consistent performance across datasets and languages—including those not encountered during the regressor’s training, such as unseen languages for the pre-trained LMs—emphasizes the model’s generalization capabilities. We also observe better performance for M2M100 compared to NLLB, which may be attributed to NLLB’s reliance on an English-centric dataset containing only seen languages, lacking examples of unseen languages for the regressor. This may indicate the importance of including instances of unseen languages in the regressor training dataset for achieving more robust predictions.
Figure 3 shows the impact of features used in ProxyLM for MT in the LOLO setting with XGBoost. Utilizing proxy models as features leads to a significant reduction in RMSE across all scenarios, showcasing their importance compared to other features. For the English-centric dataset, including language and dataset features alongside proxy models enhances performance. Dataset features alone show better improvement than language features alone, but the combination of both yields the best performance. On the other hand, for the Many-to-Many Languages dataset, the benefits of incorporating dataset and language features are less pronounced, especially for the M2M100, and there may even be a performance dip for the NLLB due to the dataset’s lack of domain shift.
Figure 2: Unseen and Cross-Dataset MT test results on English-centric dataset in average RMSE (lower is better). We only show the best-performing baseline for comparison with ProxyLM with different proxy models. “No FT" denotes “no fine-tuning". We only show M2M100 results for the Unseen setting since NLLB covers all languages in the English-centric dataset. The reported results for the Unseen setting use XGBoost, while the Cross-Dataset experiments use LGBM. Ensemble denotes combining all four proxy models. The detailed breakdown of this result with the standard deviation can be seen in the Appendix Section C.
@@ 5.1 Robustness of ProxyLM
M2M100 421 s 3.94 hrs (7.04$\times$) 1.42 hrs (7.32$\times$)
NLLB 737 s 12.08 hrs (21.57$\times$) 7.21 hrs (37.08$\times$)
@@
SMaLL-100 333 s 2.38 hrs (4.25$\times$) 1.03 hrs (5.29$\times$)
Transformer 231 s 0.56 hrs (1$\times$) 0.19 hrs (1$\times$)
Table 3: Comparison of LMs’ inference time (in seconds) and fine-tuning time (in hours) for one MT experimental run. The multiplier of fine-tuning time is relative to the Transformer model. All times were calculated using the interquartile mean to ignore outliers.
Datasets Intent Classification Slot Filling
Estimated models
Aya-23 0.68 hrs (6.13$\times$) 5.10 hrs (13.48$\times$)
LLaMA3 0.54 hrs (4.91$\times$) 4.36 hrs (11.53$\times$)
Proxy models
BLOOMZ-560M 0.12 hrs (1.06$\times$) 0.59 hrs (1.57$\times$)
SmolLM-360M 0.12 hrs (1.09$\times$) 0.40 hrs (1.05$\times$)
SmolLM-150M 0.11 hrs (1.00$\times$) 0.38 hrs (1.00$\times$)
Table 4: Comparison LMs’ overall fine-tuning and evaluation time (in hours) for one intent classification or slot filling experimental run. The multiplier of the time is relative to SmolLM-135M model as it is the smallest proxy model. All times were calculated using the interquartile mean to ignore outliers.
Figure 4: Detailed results of XGBoost with ProxyLM Ensemble on M2M100 model under the LOLO setting on MT using the English-centric dataset from Table 1. The results are grouped by (a) Joshi Class and (b) language family that follows the mapping which is provided in Appendix B.1; (c) shows the scatter plot illustrating the correlation of spBLEU scores between the ProxyLM’s prediction and estimated LM, with the light gray dashed line representing the line of equality ($y=x$) with $R^{2}=0.90$ and black dashed line representing Locally Weighted Scatterplot Smoothing (LOWESS) curve to represent the trend.
@@ 5.2 Time Efficiency
Table 3 compares the fine-tuning and inference times required for the estimated and proxy models on MT, while Table  4 compares the fine-tuning and evaluation times required for the estimated and proxy models on intent classification and slot filling. The results demonstrate that fine-tuning proxy models or direct inference from any model is remarkably faster than fine-tuning all estimated models. Table 5 further illustrates this point, showing only a minimal trade-off in the time needed to train the regressor models. This additional training time is relatively negligible, highlighting the efficiency of using proxy models.
Regressors English-centric (MT) Many-to-Many (MT) Intent and Slot
XGBoost 2.24 s 0.87 s 6.33 s
Poly2 0.07 s 0.06 s 0.09 s
Poly3 0.06 s 0.06 s 0.10 s
LGBM 90.79 s 27.51 s 118.24 s
MF N/A 141.69 s N/A
Table 5: Regressor models training time (in seconds) per one round of cross-validation with 10-folds across all setups. We combine the timing for intent classification and slot filling since they both contain the same amount of training data. All times were calculated using the interquartile mean to ignore outliers.
@@ 5.3 Performance by Language Categories
In Figure 4, we present detailed XGBoost results with ProxyLM Ensemble on the M2M100 model under the English-centric LOLO experiment, grouped by language categories. ProxyLM demonstrates relatively stable performance across languages belonging to different Joshi classes and linguistic families. Based on the Locally Weighted Scatterplot Smoothing (LOWESS)  Cleveland (1979) curve depicted in Figure 4(c), our method consistently maintains unbiased predictions for spBLEU scores below 40 across various language types. However, as the spBLEU score increases, the availability of data points diminishes, leading to our method under-predicting the performance compared to the true spBLEU score. Outliers observed in Kartvelian languages and Indo-European languages with Joshi class 3 may have contributed to this discrepancy in prediction. These observations suggest that increasing the number of data points covering higher spBLEU scores may help mitigate the bias in prediction.
@@ 6 Related Work
The prediction performance of machine learning algorithms has been mainly explored in two research directions: (1) predict the model performance during the training runtime, and (2) predict the model performance by providing extracted features from the dataset Xia et al. (2020).
Performance Prediction During the Training Runtime.
@@
Performance Prediction Using Extracted Features.
The latter aims to predict the model performance by learning a correlation between input features and the final evaluation metric. Birch et al. (2008) identify strong predictive features such as the amount of reordering, the morphological complexity of the target language, and the historical relatedness of the two languages. Xia et al. (2020) leverage extracted dataset features and typological database language representations. Ye et al. (2021) introduce the use of confidence intervals and calibration with various regressor algorithms for reliable performance prediction. Schram et al. (2023) apply Bayesian matrix factorization for performance prediction on multilingual NLP tasks. In this work, we focus on exploring the latter. Existing approaches have shown promise using linear regression and gradient-boosting trees Birch et al. (2008); Xia et al. (2020); Srinivasan et al. (2021); Ye et al. (2021). These studies have considered data size, typological features, and language similarity as factors contributing to the model performance.
Enhancing LLMs through Small Models.
Recent studies, as compiled in Chen and Varoquaux (2024), have explored leveraging smaller models to complement LLMs across various tasks. A closely related application involves using smaller models to assist with LLM inference and evaluation Kuhn et al. (2023); Manakul et al. (2023); Wang et al. (2024); Winata et al. (2024a). In contrast, our work focuses on employing proxy models within a language- and task-agnostic framework to predict LLM performance accurately. This approach offers a cost-effective alternative to fine-tuning and inference during model selection, while achieving high evaluation accuracy and robustness across extreme domain shifts. These include datasets spanning diverse domains, quality levels, and languages, with particular effectiveness demonstrated for extremely LRLs.
@@ 7 Conclusion
In this paper, we introduce ProxyLM, a novel framework designed to predict the performance of LMs by leveraging proxy models including for LRLs. By utilizing proxy models as substitutes to estimate the performance of the target model, we strategically employ smaller LMs or off-the-shelf models without additional fine-tuning. This framework is highly scalable to multiple proxy models and is task- and language-agnostic, making it applicable to a wide range of downstream NLP tasks. Our approach showcases substantial advancements in prediction accuracy compared to standard baselines and exhibits strong generalization capabilities across varied scenarios.
@@
This paper focuses on the empirical use of various proxy models without delving into the intricacies of the proxy model selection. Specifically, we do not investigate methods for determining which proxy models are the most effective across different contexts without relying on empirical experimentation. This limitation highlights a significant avenue for future research, which could involve a more systematic approach to identifying the most effective proxy models given estimated LMs and NLP tasks. Nevertheless, the use of proxy models consistently outperforms the absence of such models, demonstrating their importance in performance prediction.
Alternatively, developing robust methodologies for collecting and analyzing relevant past performance records could provide invaluable insights that enhance the generalization and accuracy of our predictive framework. Some performance data may offer greater information gain than others, potentially minimizing the number of performance records required to achieve a more robust and accurate predictor. By establishing a clearer understanding of which performance records yield the most informative insights, we can optimize our approach and improve our overall predictive capabilities. Future research in this area may further explore various data collection strategies and analytical techniques to develop a more comprehensive framework for selecting and utilizing proxy models effectively.
Ethical Considerations
We are committed to conducting our evaluations with the utmost standards of transparency and fairness. This commitment involves applying rigorous methodologies that ensure equitable and unbiased assessment processes.